OpenStudy (mommyo61):

please help, thank you so much. the barge travels upstream 75.5 miles to the factory and then back to the warehouse.the barge travels upstream at a rate of 5 miles per hour slower than it travels downstream.the barge total time is 15 hours.what is the rate the barge travels downstream?

OpenStudy (sshayer):

let speed upstream=x mi/hr then downstream speed=x+5 m/hr distance =75.5 miles \[\frac{ 75.5 }{ x }+\frac{ 75.5 }{ x+5 }=15\] calculate x

OpenStudy (mommyo61):

thank you so much. I tried everything but that.

OpenStudy (sshayer):

yw.